• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name TO

    Log entry time 13:22:33 on June11,2002

    Entry number 85887

    keyword=target H2 temperature

    At the begining of run #1443, the temperature of the target was around 19.5 K, and both temperature and pressure were not stable (this was due to the difficulty MCC have had to deliver 90 muA beam).
    But a few time later (after the first 50000 evts ?) everything was good.